1. What are the general equations of an ellipse?


2. What is the general equation of a circle?


3. What is the general equation of a straight line?


4. What is the general equation of a parabola that opens upward with V(-1,2)?


5. What is the equation of a hyperbola with major axis parallel to the y – axis?


1
Expert's answer
2021-01-28T05:40:04-0500

Solution 1:

General equations of the ellipse:

Case I: Major axis is y-axis (vertical axis)

"\dfrac{(x-h)^2}{b^2}+\dfrac{(y-k)^2}{a^2}=1" , where "a>b" and (h, k) is centre.

Case II: Major axis is x-axis (horizontal axis)

"\dfrac{(x-h)^2}{a^2}+\dfrac{(y-k)^2}{b^2}=1" , where "a>b" and (h, k) is centre.

Solution 2:

General equation of the circle:

"(x-h)^2+(y-k)^2=r^2" , where r is radius and (h, k) is centre.

Solution 3:

General equation of straight line:

"y=mx+c" , where m is slope or gradient and c is y-intercept

Solution 4:

The parabola opens upward means parabola's axis is parallel to y-axis.

And we know general equation of parabola parallel to y-axiswith vertex (h, k):

"y-k=a(x-h)^2"

Given, (h, k) = (-1, 2)

We get, "y-2=a(x+1)^2"

Required general equation is "y-2=a(x+1)^2"

Solution 5:

Equation of hyperbola with major axis parallel to y-axis:

"\dfrac{(y-k)^2}{a^2}-\dfrac{(x-h)^2}{b^2}=1" , where "a>b" and (h, k) is vertex.